OpenStudy (anonymous):

Reverse the order of integration: S(0 to 1) S(0 to sqrt(x)) (x^3 +1)^(1/3) dxdy....Thank you for all the help!

OpenStudy (anonymous):

\[\int\limits_{0}^{1}\int\limits_{0}^{\sqrt{y}}\sqrt{x^{3}+1}dxdy\] This is correct =)